# Bi-Lipschitz embedding metric triangles in the plane

### v1

@@ 3 Proof of optimal distortion bound
Observe that $\bar{x}$ and $\bar{y}$ belong to the same leg of the embedded tripod $Y$. First we show the upper bound. Note that $d(x,y)\geq|t-s|$, since by 2.7 the map $\varphi\circ P$ is $1$-Lipschitz. It follows that $d(x,y)\geq\max\{a,b,|t-s|\}$. Let $e=|t-s|$. A computation shows that
@@
The function $M_{1}(b,e)$ is maximized on the set $\{(b,e):0\leq b,e\leq 1\}$ by taking $b=e=1$, which gives a value of $4$. Since this statement is plausible enough (for example, by plotting the function), and checking it is unrelated to the rest of the proof, we defer the verification of this claim to the Appendix. See 6.1 for the details. We do the same with similar inequalities arising in this section.
@@
On the set $\{(b,e):0\leq b\leq 1,1\leq e<\infty\}$, the function $e^{-2}M_{1}(b,e)$ is maximized by again taking $b=e=1$, which gives a value of $4$. See 6.2. This establishes the upper bound.
@@
Assume without loss of generality that $a=1$. Set $e=s-t$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
In 6.3, we show that $M_{2}(b,e)$ is bounded on the set $\{(b,e):b\geq 0\}$ by $4$.
@@
By 6.4, the function $\widetilde{M}_{2}(b,e)$ is bounded on the set $\{(b,e):b\geq 0,-b\leq e\leq 0\}$ by $13/3$. This covers the case where $b\geq t-s$.
Assume next that $w\in I_{1}$ with $d(p,w)\leq d(p,x)$. Since $w\in I_{1}$, we can follow the previous argument with roles of $x$ and $y$ reversed. Since $d(p,w)\leq d(p,x)$, we get the inequality $d(x,y)\leq 2b+t-s$ analogously to the first alternative above, which is handled by 6.3.
@@
If $d_{1}\geq 0$, we use the first inequality and the relation $d(p,q)=d(p,o_{3})+d(o_{2},q)$ to conclude that $d_{1}\leq b-t$. If $d_{1}\leq 0$, we use the second inequality and the relation $d(p,r)=d(p,o_{3})+d(o_{2},r)$ to conclude that $-d_{1}\leq b-t$. Together, these show that $|d_{1}|\leq b-t$. Similarly, we have $|d_{2}|\leq a-s$.
@@
Note as well that $s\leq a=1$ and $t\leq b\leq 1$. Thus we must maximize the function $M_{3}(b,s,t)$ on the set $0\leq b,s,t\leq 1$. We show in 6.5 that the maximum value is $16/3$, which is attained when $b=1$ and $s=t=0$. ∎
The remaining two cases are handled similarly to 3.3, and so our proofs omit some details.
@@
We begin with the upper bound. Observe as before that $d(x,y)\geq\max\{a,b,s+t\}$. On the other hand, a computation shows that
@@
Over the region $\{(b,s,t):0\leq b,s,t\leq 1,s+t\leq 1\}$, the expression on the right of the previous inequality is maximized by taking $s=0$, $b=t=1$ by 6.6. This gives the desired upper bound $71/4+3/4\cdot 3^{2}=$ 7.
@@
Over the region $\{(b,s,t):0\leq s,t\leq 1,s+t\geq 1,0\leq b\leq s+t\}$, the expression on the right of the previous inequality is again maximized by taking $s=0$, $b=t=1$ by 6.7.
@@
Over the region $\{(b,s,t):0\leq s,t\leq 1,b\geq 1,s+t\leq b\}$, the expression on the right of the previous inequality is again maximized by taking $s=0$, $b=t=1$ by 6.8.
@@
$M_{5}(b,s,t)=\frac{(\max\{2,2s\}+s+t)^{2}}{\frac{1}{4}\left(1-b+2s+t\right)^{2}+\frac{3}{4}\left(1+t+b\right)^{2}}.$
By 6.9, the function $M_{5}(b,s,t)$ is bounded by $4$ on the set $\{(b,s,t):0\leq b,s,t,1\geq s\}$.
Case 3. In the final case, we assume that $z\in I_{2}$ and $w\in I_{2}$. Proceeding similarly to Case 2 of 3.3, we obtain the relation $d(x,y)\leq 2a+2b-s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}\leq M_{6}(b,s,t)$, where
@@
Note that the numerator of this formula uses the weaker bound $2+2b+2t$ instead of $2+2b-s+t$, but this turns out to simplify the calculations. By 6.10, this is bounded on the set $\{(b,s,t):0\leq b,s,t\}$ by $16/3$. ∎
@@
Suppose that $x\in I_{1}^{p}$ and $y\in I_{3}^{r}$. Then
@@
We prove the upper bound. This is similar to the proof of 3.4. As before, we have the lower bound $d(x,y)\geq\max\{a,b,s+t\}$. However, we have the new computation
@@
Over the region $\{(b,s,t):0\leq b,s,t\leq 1,s+t\leq 1\}$, $M_{7}(b,s,t)$ is maximized by taking $s=0$, $t=1$, $b=1$ or $s=1$, $t=0$, $b=1$. By 6.11, this gives the value $71/4+3/4\cdot 3^{2}=$ 7, which gives the desired bound.
@@
Over the region $\{(b,s,t):0\leq b,s,t,1\leq s+t\}$, the function $M_{7}(b,s,t)/(s+t)^{2}$ is maximized by taking $s=0$, $t=1$, $b=1$ or $s=1$, $t=0$, $b=1$, which by 6.12 again gives the desired bound.
Next, we show the lower bound. Assume first that $z\in I_{3}$. Then we can show similarly to in 3.4 that $d(y,z)\leq a+t-s$. Actually, for the remainder of the proof it is slightly more convenient to use the weaker bound $d(y,z)\leq a+t+s$. From this we conclude that $d(x,y)\leq 2a+s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
where $b,s,t\geq 0$. By 6.13, this function is bounded on the given region by $4$.
@@
Finally, assume that both $z$ and $w$ are in $I_{2}$. We can show similarly to Case 2 in 3.3 that $d(w,z)\leq a+b+s+t$. From this we conclude that $d(x,y)\leq 2a+2b+s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
where $b,s,t\geq 0$. As in Case 3 of 3.4, we use a weaker bound in the numerator in order to simplify the calculations. By 6.14, this function is bounded above by $16/3$. ∎
4 Examples
We give a detailed analysis of the examples introduced in 2.5: the intrinsic circle, the three-petal rose, tripods and the twisted heart. Theses are denoted respectively by $S$, $R$, $T$ and $H$.
4.1 The intrinsic circle
This example represents the “fattest” possible metric triangle.
The obvious embedding from $S$ into $\mathbb{R}^{2}$ is the identity map, which has distortion $\pi/2\approx 1.5708$. Interestingly, the identity map is not distortion minimizing. To improve this, for each $s\in[0,\pi/3)$, we define the map $F_{s}\colon S\to\mathbb{R}^{2}$ by the formula
$F_{s}(\theta)=\left(\frac{3\pi\cos(\tau(\theta))-\sqrt{3}\pi\sin(s)}{3\pi-9s},\frac{\pi(-\sin(s)+\sin(\tau(\theta)))}{\pi-3s}\right)$
for $\theta\in[0,2\pi/3)$. Here, $\tau(\theta)=s+(1-3s/\pi)\theta$. We extend the definition to all $\theta\in[0,2\pi)$ by symmetry. Namely, let $A\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ be the linear map representing rotation by $\pi/3$. For $\theta\in[2\pi/3,2\pi)$, we define $F_{s}(\theta)=A\circ F_{s}(\theta-2\pi/3)$. The map $F_{s}$ takes $S$ to a “fattened triangle” as shown in Figure 5 (with the value $s_{0}$ found below). Note that $F_{0}$ is the identity map and that the limit of $F_{s}$ as $s\to\pi/3$ maps $S$ onto an equilateral triangle. The intuition behind the family $\{F_{s}\}$ is that, beginning from the identity map, one can lower the maximal distortion by adding corners, which increases the distortion on small scales (where the identity map is close to an isometry) but lowers distortion between antipodal points (where the maximal distortion is attained).
$p$$q$$r$
Figure 5: The image of $F_{s_{0}}$, which embeds the intrinsic circle in $\mathbb{R}^{2}$
The map $F_{s}$ is length-preserving, and in particular is $1$-Lipschitz but not $L$-Lipschitz for any $L<1$. Thus the distortion of $F_{s}$ is given by finding two points $x,y\in S$ that maximize $d(x,y)/\|F_{s}(x)-F_{s}(y)\|$. A calculation shows that the distortion of $F_{s}$ is attained by the pair of antipodal points $F_{s}(0)$ and $F_{s}(\pi)$ if $s$ is not too large. For such $s$ we calculate the distortion of $F_{s}$ to be
$\operatorname{lip}(F_{s})=\frac{\pi-3s}{1+\cos(s)-\sqrt{3}\sin(s)}.$
[…]
Question 4.1.
What is $\operatorname{lip}(S,\mathbb{R}^{2})$, and what embedding attains it?
In the other direction, an easy argument shows that $\operatorname{lip}(S,\mathbb{R}^{2})\geq\sqrt{2}$. Let $x,y,z,w\in S$ be four evenly spaced points, so that $d(x,y)=d(y,z)=d(z,w)=d(w,x)=\pi/2$ and $d(x,z)=d(y,w)=\pi$. Let $f\colon S\to\mathbb{R}^{2}$ be an arbitrary embedding. By rescaling, we may assume that $f$ is $1$-Lipschitz. A consequence of Euler’s quadrilateral theorem is the inequality
$\|f(x)-f(z)\|^{2}+\|f(y)-f(w)\|^{2}\leq\|f(x)-f(y)\|^{2}+\|f(y)-f(z)\|^{2}+\|f(z)-f(w)\|^{2}+\|f(w)-f(x)\|^{2}.$
The right-hand side of this inequality is bounded by $\pi^{2}$. This implies that either $\|f(x)-f(z)\|\leq\pi/\sqrt{2}$ or $\|f(y)-f(w)\|\leq\pi/\sqrt{2}$. In the first case we have
$\operatorname{lip}(f)\geq\frac{d(x,z)}{\|f(x)-f(z)\|}\geq\sqrt{2}.$
The second case likewise gives $\operatorname{lip}(f)\geq\sqrt{2}$.
4.2 The three-petal rose
@@
Choose a value $\varepsilon\in(0,1/(2\operatorname{lip}(G)))$. Observe that $d(G(y_{i}),0)\geq\varepsilon$ for each $i$. In particular, since each set $G(S_{i}^{j})$ is connected, we can find a point $x_{i}^{j}\in S_{i}^{j}$ such that $d(f(x_{i}^{j}),0)=\varepsilon$.
@@
and hence that the distortion of $G$ is at least $22\varepsilon/\varepsilon=$ 2. ∎
@@
4.3 Tripods
As we have seen, tripods play a key role in the proof of 1.1. The class of tripods represents the “thinnest” possible metric triangles. We verify the natural expectation that the embedding $F$ from the previous section is distortion minimizing for this class of triangles.
@@ 4.3 Tripods
The proof that no embedding can have smaller distortion is similar to the proof of 4.2 but with three arcs emanating from the center point $o$ rather than six. We omit the details. ∎
4.4 The twisted heart
This example shows the sharpness of 1.1. Observe first that $d(b,e)=d(c,a)=1$ and $d(a,b)=4$. On the other hand, we compute that $F(a)=(\frac{1}{2},\frac{\sqrt{3}}{2})$, $F(b)=(\frac{1}{2},-\frac{\sqrt{3}}{2})$ and $F(c)=(-\frac{3}{2},-\frac{\sqrt{3}}{2})$. Then
@@ 4.4 The twisted heart
5 Metric polygons
In this section, we prove 1.3, which shows that the embedding result for metric triangles does not extend to the case of metric quadrilaterals.
@@ 5 Metric polygons
1.3 is a consequence of the following proposition.
@@
$p_{1}^{\prime}$$p_{2}^{\prime}$$p_{3}^{\prime}$$p_{4}^{\prime}$$m_{1}^{\prime}$$m_{2}^{\prime}$
@@
Figure 6: The metric polygons in Section 5.
@@
Assume that there is an $L$-bi-Lipschitz embedding $f\colon Q\to\mathbb{R}^{2}$. Given a point $x\in Q$, we write $x^{\prime}$ in place of $f(x)$.
Observe that $f(Q)$ is a simple closed curve that, by the Jordan curve theorem, separates the plane and has two complementary components.
@@
Let $J_{1}$ be the straight line segment connecting $p_{1}^{\prime}$ to $p_{3}^{\prime}$, which has length at most $L\varepsilon$. We can find a maximal closed subinterval $\widetilde{J}_{1}$ with interior contained in $\mathbb{R}^{2}\setminus f(Q)$ such that one endpoint $j_{1}^{1}\in\widetilde{J}_{1}$ belongs to $f(I_{1}\cup I_{4})$ and the other endpoint $j_{1}^{2}\in\widetilde{J}_{1}$ belongs to $f(I_{2}\cup I_{3})$. Let $\widetilde{J}_{1}^{1}$ be the minimal subarc of $f(I_{1}\cup I_{4})$ connecting $p_{1}^{\prime}$ to $j_{1}^{1}$, and let $\widetilde{J}_{1}^{2}$ be the minimal subarc of $f(I_{2}\cup I_{3})$ connecting $j_{1}^{2}$ to $p_{3}^{\prime}$. Observe that $\widetilde{J}_{1}^{1}$ and $\widetilde{J}_{1}^{2}$ each have length at most $L^{3}\varepsilon$, since
$\ell(\widetilde{J}_{1}^{1})\leq L\ell(f^{-1}(\widetilde{J}_{1}^{1}))=Ld(p_{1},f^{-1}(j_{1}^{1}))\leq L^{2}\|p_{1}^{\prime}-j_{1}^{1}\|\leq L^{3}\varepsilon,$
and similarly for $\widetilde{J}_{1}^{2}$. Then the concatenation $K_{1}=\widetilde{J}_{1}^{1}*\widetilde{J}_{1}*\widetilde{J}_{1}^{3}$ is a simple curve having length at most $3L^{3}\varepsilon$.
We define $K_{2}$ in the same way as a simple curve with length at most $3L^{3}\varepsilon$ connecting $p_{2}^{\prime}$ to $p_{4}^{\prime}$.
If $K_{1}$ and $K_{2}$ were to have a point of intersection, it would follow that $\|p_{1}^{\prime}-p_{2}^{\prime}\|\leq 6L^{3}\varepsilon$, and hence that $d(p_{1},p_{2})\leq 6L^{4}\varepsilon<1$. Since $d(p_{1},p_{2})=1$, we conclude that $K_{1}$ and $K_{2}$ do not have a point of intersection. In particular, it follows that $\widetilde{J}_{1}$ and $\widetilde{J}_{2}$ lie in different complementary components of $f(Q)$.
@@
Let $\widetilde{S}$ be the union of $K_{1}$, $K_{2}$, $f(I_{2})$ and $f(I_{4})$. We observe that $\widetilde{S}$ separates $m_{1}^{\prime}$ and $m_{3}^{\prime}$. Then $\widetilde{S}$ contains a simple closed curve $S$ also separating $m_{1}^{\prime}$ and $m_{3}^{\prime}$ (see Theorem IV.6.7 in [12]). Now $\|m_{1}^{\prime}-m_{3}^{\prime}\|\leq L\varepsilon$, which implies that $m_{1}^{\prime}$ and $m_{3}^{\prime}$ lie at distance at most $L\varepsilon<(6L^{3})^{-1}\leq(6L)^{-1}$ from $S$.
However, the distance from $m_{1}^{\prime}$ to $K_{1}\cup K_{2}$ is at least
$\frac{1}{2L}-3L^{2}\varepsilon\geq\frac{1}{2L}-\frac{1}{6L^{4}}\geq\frac{1}{4L}.$
Likewise, the distance from $m_{1}^{\prime}$ to $f(I_{2})\cup f(I_{4})$ is at least $(2L)^{-1}$. Thus $\operatorname{dist}(m_{1}^{\prime},S)\geq(4L)^{-1}$, which gives a contradiction. ∎
@@
We provide the details for $K_{5}$. Denote its topological vertices in cyclic order by $p_{1},p_{2},\ldots,p_{5}$. Denote the topological edge from $p_{i}$ to $p_{j}$ by $I_{ij}$. We make $K_{5}$ into a metric space by assigning each edge length $1$ and taking the length metric. For each $i=1,\ldots,5$, let $m_{i}$ be the midpoint of $I_{i(i+1)}$ (with the usual identification of $I_{56}$ with $I_{51}$). The space $K_{5}$ can be viewed as a metric pentagon by taking $m_{1},m_{3},m_{5},m_{2},m_{4}$ as its vertices (as a metric polygon) in cyclic order. The edges (as a metric polygon) are $[m_{1}p_{1}]*I_{14}*[p_{4}m_{3}]$, $[m_{3}p_{3}]*I_{31}*[p_{1}m_{5}]$, $[m_{5}p_{5}]*I_{53}*[p_{3}m_{2}]$, $[m_{2}p_{2}]*I_{25}*[p_{5}m_{4}]$, $[m_{4}p_{4}]*I_{42}*[p_{2}m_{1}]$. Each of these edges has length $2$.

### v2

@@ 3. Proof of optimal distortion bound
Observe that $\bar{x}$ and $\bar{y}$ belong to the same leg of the embedded tripod $Y$. First we show the upper bound. Note that $d(x,y)\geq|t-s|$, since by ˜2.7 the map $\varphi\circ P$ is $1$-Lipschitz. It follows that $d(x,y)\geq\max\{a,b,|t-s|\}$. Let e=|ts|e=|t-s|. Observe that $F(x)=(s+\frac{1}{2}a,\frac{\sqrt{3}}{2}a)$ and $F(y)=(t+\frac{1}{2}b,-\frac{\sqrt{3}}{2}b)$, from which we calculate
@@
The function $M_{1}(b,e)$ is maximized on the set $\{(b,e):0\leq b,e\leq 1\}$ by taking $b=e=1$, which gives a value of $4$. Since this statement is plausible enough (for example, by plotting the function), and checking it is unrelated to the rest of the proof, we defer the verification of this claim to the Appendix. See ˜6.1 for the details. We do the same with similar inequalities arising in this section.
@@
On the set $\{(b,e):0\leq b\leq 1,1\leq e<\infty\}$, the function $e^{-2}M_{1}(b,e)$ is maximized by again taking $b=e=1$, which gives a value of $4$. See ˜6.2. This establishes the upper bound.
@@
Assume without loss of generality that $a=1$. Set e=ste=s-t. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
In ˜6.3, we show that $M_{2}(b,e)$ is bounded on the set $\{(b,e):b\geq 0\}$ by $4$.
@@
By ˜6.4, the function $\widetilde{M}_{2}(b,e)$ is bounded on the set $\{(b,e):b\geq 0,-b\leq e\leq 0\}$ by $13/3$. This covers the case where $b\geq t-s$.
Assume next that $w\in I_{1}$ with $d(p,w)\leq d(p,x)$. Since $w\in I_{1}$, we can follow the previous argument with roles of $x$ and $y$ reversed. Since $d(p,w)\leq d(p,x)$, we get the inequality $d(x,y)\leq 2b+t-s$ analogously to the first alternative above, which is handled by ˜6.3.
@@
If $d_{1}\geq 0$, we use the first inequality and the relation $d(p,q)=d(p,o_{3})+d(o_{2},q)$ to conclude that $d_{1}\leq b-t$. If $d_{1}\leq 0$, we use the second inequality and the relation $d(p,r)=d(p,o_{3})+d(o_{2},r)$ to conclude that $d_{1}\leq b-t$. Together, these show that $|d_{1}|\leq b-t$. Similarly, we have $|d_{2}|\leq a-s$.
@@
Note as well that $s\leq a=1$ and $t\leq b\leq 1$. Thus we must maximize the function $M_{3}(b,s,t)$ on the set $0\leq b,s,t\leq 1$. We show in ˜6.5 that the maximum value is $16/3$, which is attained when $b=1$ and $s=t=0$. ∎
The remaining two cases are handled similarly to ˜3.3, and so our proofs omit some details.
@@
We begin with the upper bound. Observe as before that $d(x,y)\geq\max\{a,b,s+t\}$ and that $F(x)=(s+\frac{1}{2}a,\frac{\sqrt{3}}{2}a)$. On the other hand, now $F(y)=(\frac{1}{2}(-t+b),-\frac{\sqrt{3}}{2}(t+b))$, and we obtain
@@
Over the region $\{(b,s,t):0\leq b,s,t\leq 1,s+t\leq 1\}$, the expression on the right of the previous inequality is maximized by taking $s=0$, $b=t=1$ by ˜6.6. This gives the desired upper bound $1/4+3/4\cdot 3^{2}=7$.
@@
Over the region $\{(b,s,t):0\leq s,t\leq 1,s+t\geq 1,0\leq b\leq s+t\}$, the expression on the right of the previous inequality is again maximized by taking $s=0$, $b=t=1$ by ˜6.7.
@@
Over the region $\{(b,s,t):0\leq s,t\leq 1,b\geq 1,s+t\leq b\}$, the expression on the right of the previous inequality is again maximized by taking $s=0$, $b=t=1$ by ˜6.8.
@@
$M_{5}(b,s,t)=\frac{(\max\{2,2b\}+s+t)^{2}}{\frac{1}{4}\left(1-b+2s+t\right)^{2}+\frac{3}{4}\left(1+t+b\right)^{2}}.$
By ˜6.9, the function $M_{5}(b,s,t)$ is bounded by $13/3$ on the set $\{(b,s,t):0\leq b,s,t,1\geq s\}$.
Case 3. In the final case, we assume that $z\in I_{2}$ and $w\in I_{2}$. Proceeding similarly to Case 2 of ˜3.3, we obtain the relation $d(x,y)\leq 2a+2b-s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}\leq M_{6}(b,s,t)$, where
@@
Note that the numerator of this formula uses the weaker bound $2+2b+2t$ instead of 2+2bs+t2+2b-s+t, but this turns out to simplify the calculations. By ˜6.10, this is bounded on the set $\{(b,s,t):0\leq b,s,t\}$ by $16/3$. ∎
@@
Suppose that $x\in I_{1}^{q}$ and $y\in I_{3}^{r}$. Then
@@
We prove the upper bound. This is similar to the proof of ˜3.4. As before, we have the lower bound $d(x,y)\geq\max\{a,b,s+t\}$. However, in this case we have $F(x)=(\frac{1}{2}(-s+a),\frac{\sqrt{3}}{2}(s+a))$ $F(y)=(\frac{1}{2}(-t+b),-\frac{\sqrt{3}}{2}(t+b))$, and so
@@
Over the region $\{(b,s,t):0\leq b,s,t\leq 1,s+t\leq 1\}$, $M_{7}(b,s,t)$ is maximized by taking $s=0$, $t=1$, $b=1$ or $s=1$, $t=0$, $b=1$. By ˜6.11, this gives the value $1/4+3/4\cdot 3^{2}=7$, which gives the desired bound.
@@
Over the region $\{(b,s,t):0\leq b,s,t,1\leq s+t\}$, the function $M_{7}(b,s,t)/(s+t)^{2}$ is maximized by taking $s=0$, $t=1$, $b=1$ or $s=1$, $t=0$, $b=1$, which by ˜6.12 again gives the desired bound.
Next, we show the lower bound. Assume first that $z\in I_{3}$. Then we can show similarly to in ˜3.4 that $d(y,z)\leq a+t-s$. Actually, for the remainder of the proof it is slightly more convenient to use the weaker bound $d(y,z)\leq a+t+s$. From this we conclude that $d(x,y)\leq 2a+s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
where $b,s,t\geq 0$. By ˜6.13, this function is bounded on the given region by $4$.
@@
Finally, assume that both $z$ and $w$ are in $I_{2}$. We can show similarly to Case 2 in ˜3.3 that $d(w,z)\leq a+b+s+t$. From this we conclude that $d(x,y)\leq 2a+2b+s+t$. Assume without loss of generality that $a=1$. Then $d(x,y)^{2}/\|F(x)-F(y)\|^{2}$ is bounded by the function
@@
where $b,s,t\geq 0$. As in Case 3 of ˜3.4, we use a weaker bound in the numerator in order to simplify the calculations. By ˜6.14, this function is bounded above by $16/3$. ∎
4. Examples
We give a detailed analysis of the examples introduced in ˜2.5: the intrinsic circle, the three-petal rose, tripods and the twisted heart. Theses are denoted respectively by $S$, $R$, $T$ and $H$.
4.1. The intrinsic circle
This example represents the “fattest” possible metric triangle. The obvious embedding from $S$ into $\mathbb{R}^{2}$ is the identity map, which has distortion $\pi/2$. This embedding is indeed distortion-minimizing; this follows from a result of Linial–Magen [9, Claim 2.1] on embedding $n$-cycles into Euclidean space. See also Theorem 35 in [3].
4.2. The three-petal rose
@@
Choose a value $\varepsilon\in(0,1/(2\operatorname{lip}(G)))$. Observe that $\|G(y_{i})\|\geq\varepsilon$ for each $i$. In particular, since each set $G(S_{i}^{j})$ is connected, we can find a point $x_{i}^{j}\in S_{i}^{j}$ such that $\|G(x_{i}^{j})\|=\varepsilon$.
@@
and hence that the distortion of $G$ is at least $2\varepsilon/\varepsilon=2$. ∎
@@
4.3. Tripods
As we have seen, tripods play a key role in the proof of ˜1.1. The class of tripods represents the “thinnest” possible metric triangles. We verify the natural expectation that the embedding $F$ from the previous section is distortion minimizing for this class of triangles. Note in this case that $d(x,\widehat{I}_{j})=0$ for all $x\in I_{j}$ ($j=1,2,3$), and so $F(x)=\bar{x}$ for all points $x$.
@@ 4.3. Tripods
The proof that no embedding can have smaller distortion is similar to the proof of ˜4.1 but with three arcs emanating from the center point $o$ rather than six. We omit the details. ∎
4.4. The twisted heart
This example shows the sharpness of ˜1.1. Observe first that $d(b,e)=d(c,a)=1$ and $d(a,b)=4$. Moreover, $\bar{a}=\bar{b}=(0,0)$ and $\bar{c}=(-\frac{1}{2},\frac{\sqrt{3}}{2})$. We now compute that $F(a)=\bar{a}+v_{1}=(\frac{1}{2},\frac{\sqrt{3}}{2})$, $F(b)=\bar{b}+\bar{v}_{3}=(\frac{1}{2},-\frac{\sqrt{3}}{2})$ and $F(c)=\bar{c}+v_{2}=(-\frac{3}{2},-\frac{\sqrt{3}}{2})$. This yields the equalities
@@ 4.4. The twisted heart
5. Metric polygons
In this section, we prove ˜1.3, which shows that the embedding result for metric triangles does not extend to the case of metric quadrilaterals.
@@ 5. Metric polygons
˜1.3 is a consequence of the following proposition.
@@
$j_{1}$$j_{2}$$j_{3}$$j_{4}$$m_{1}^{\prime}$$m_{3}^{\prime}$
@@
Figure 5. The metric polygons in Section˜5.
@@
Assume that there is an $L$-bi-Lipschitz embedding $f\colon Q\to\mathbb{R}^{2}$. Given a point $x\in Q$, we write $x^{\prime}$ in place of $f(x)$. Observe that $f(Q)$ is a simple closed curve that, by the Jordan curve theorem, separates the plane and has two complementary components.
@@
Let $\widetilde{J}_{1}$ be the straight line segment connecting $p_{1}^{\prime}$ to $p_{3}^{\prime}$, which has length at most $L\varepsilon$. We can find a maximal closed subinterval $J_{1}$ with interior contained in $\mathbb{R}^{2}\setminus f(Q)$ such that one endpoint $j_{1}\in\widetilde{J}_{1}$ belongs to $f(I_{1}\cup I_{4})$ and the other endpoint $j_{3}\in\widetilde{J}_{1}$ belongs to $f(I_{2}\cup I_{3})$. Observe that necessarily $\|p_{1}^{\prime}-j_{1}\|<L\varepsilon$ and $\|p_{3}^{\prime}-j_{3}\|<L\varepsilon$. We define $J_{2}$ and points $j_{2}$ and $j_{4}$ likewise, with $\|p_{2}^{\prime}-j_{2}\|<L\varepsilon$ and $\|p_{4}^{\prime}-j_{4}\|<L\varepsilon$. We further observe that $J_{1}$ and $J_{2}$ do not intersect; if there were a point $z\in J_{1}\cup J_{2}$, then we would have $\|p_{1}^{\prime}-p_{2}^{\prime}\|\leq\|p_{1}^{\prime}-z\|+\|z-p_{2}^{\prime}\|\leq 2L\varepsilon$, which contradicts the statement that $\|p_{1}^{\prime}-p_{2}^{\prime}\|\geq 1/L$. Thus $J_{1}$ and $J_{2}$ lie in different complementary components of $f(Q)$.
We claim that the points $j_{1}$, $m_{1}^{\prime}$, $j_{2}$, $j_{3}$, $m_{3}^{\prime}$, $j_{4}$ are arranged cyclically along the simple closed curve $f(Q)$ in this order. To see this, let $H_{i}$ denote the shorter subarc of $f(Q)$ from $p_{i}^{\prime}$ to $j_{i}$ for each $1\leq i\leq 4$. Each $H_{i}$ has length at most $L^{3}\varepsilon$, since
$\ell(H_{i})\leq L\ell(f^{-1}(H_{i}))=Ld(p_{1},f^{-1}(j_{1}))\leq L^{2}\|p_{1}^{\prime}-j_{1}\|\leq L^{3}\varepsilon.$
On the other hand, $\|m_{1}^{\prime}-p_{1}^{\prime}\|\geq 1/(2L)>L^{3}\varepsilon$, which implies that $m_{1}^{\prime}$ cannot lie on the subarc $H_{1}$. Similar relations hold for other pairs of points. We conclude that the points $j_{1}$, $m_{1}^{\prime}$, $j_{2}$, $j_{3}$, $m_{3}^{\prime}$, $j_{4}$ are ordered the same as $p_{1}^{\prime}$, $m_{1}^{\prime}$, $p_{2}^{\prime}$, $p_{3}^{\prime}$, $m_{3}^{\prime}$, $p_{4}^{\prime}$, which establishes the claim.
@@
Now let $K_{1}$ be the subarc of $f(Q)$ from $j_{2}$ to $j_{3}$ and $K_{2}$ be the subarc of $f(Q)$ from $j_{4}$ to $j_{1}$ given by the previous ordering, and let $S$ be the concatenation of $K_{1}$, $J_{1}$, $K_{2}$ and $J_{2}$. Then $S$ is a simple closed curve for which $m_{1}^{\prime}$ and $m_{3}^{\prime}$ lie in different complementary components. See Figure˜5(b). Now $\|m_{1}^{\prime}-m_{3}^{\prime}\|\leq L\varepsilon$, which implies that $m_{1}^{\prime}$ and $m_{3}^{\prime}$ lie at distance at most $L\varepsilon<(6L^{3})^{-1}\leq(6L)^{-1}$ from $S$.
On the other hand, the distance from $m_{1}^{\prime}$ to $K_{1}$ is at least
$\text{dist}_{\text{Euc}}(m_{1}^{\prime},f(I_{2}))-\max\{\ell(H_{2}),\ell(H_{3})\}\geq\frac{1}{2L}-L^{3}\varepsilon\geq\frac{1}{2L}-\frac{1}{6L}\geq\frac{1}{3L}.$
Likewise, the distance from $m_{1}^{\prime}$ to $J_{1}$ is at least
$\|m_{1}^{\prime}-p_{1}^{\prime}\|-\ell(\widetilde{J}_{1})\geq\frac{1}{2L}-L\varepsilon\geq\frac{1}{3L}.$
The same bound holds for $K_{2}$ and $J_{2}$, and so $d_{\text{Euc}}(m_{1}^{\prime},S)\geq(3L)^{-1}$. This gives a contradiction, and the proposition follows. ∎
@@
We provide the details for $K_{5}$. Denote its topological vertices in cyclic order by $p_{1},p_{2},\ldots,p_{5}$. Denote the topological edge from $p_{i}$ to $p_{j}$ by $I_{ij}$. We make $K_{5}$ into a metric space by assigning each edge length $1$ and taking the length metric. For each $i=1,\ldots,5$, let $m_{i}$ be the midpoint of $I_{i(i+1)}$ (with the usual identification of $I_{56}$ with $I_{51}$). The space $K_{5}$ can be viewed as a metric pentagon by taking $m_{1},m_{3},m_{5},m_{2},m_{4}$ as its vertices (as a metric polygon) in cyclic order. The edges (as a metric polygon) are $m_{1}p_{1}]*I_{14}*[p_{4}m_{3}]$, $m_{3}p_{3}]*I_{31}*[p_{1}m_{5}]$, $m_{5}p_{5}]*I_{53}*[p_{3}m_{2}]$, $m_{2}p_{2}]*I_{25}*[p_{5}m_{4}]$, $m_{4}p_{4}]*I_{42}*[p_{2}m_{1}]$. Each of these edges has length $2$.